# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v2

@@ Abstract
Agentic AI systems, built upon large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligence, autonomy, collaboration, and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based Agentic Multi-Agent Systems (AMAS). We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents. We then adapt and extend the AI TRiSM framework for Agentic AI, structured around four key pillars: Governance, Explainability, ModelOps, and Privacy/Security , each contextualized to the challenges of multi-agent LLM systems. A novel risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation. To support practical assessment in Agentic AI works, we introduce two novel metrics: the Component Synergy Score (CSS), which quantifies the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which evaluates the efficiency of tool use within agent workflows. We further discuss strategies for improving explainability in Agentic AI , as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance. The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, outlining critical directions to align emerging systems with TRiSM principles for safe, transparent, and accountable operation.
@@ 1.1 Background and Motivation
The global market for Artificial Intelligence (AI) agents is projected to grow from \$5.4 billion in 2024 to \$7.6 billion in 2025  [1]. By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [2]. An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals  [3], has undergone a fundamental transformation. While early agents were task-specific and deterministic, modern systems have evolved significantly ,increasing operational autonomy by over 70% on benchmark tasks [4]. These Agentic AI systems now integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [5, 6].
This rapid evolution presents new opportunities but also surfaces significant challenges in coordination, control, transparency, and risk assessment across heterogeneous agents. In particular, the shift toward Agentic AI architectures amplifies and evolves longstanding concerns around safety, fairness, interpretability, and systemic harms. While such issues have been central to machine learning (ML) systems broadly, they become particularly critical in Agentic AI, where autonomous agents operate in high-stakes domains like healthcare, law, and public services, areas where their decisions can carry profound and lasting societal consequences.
From isolated agents to Agentic AI The autonomous software agents date back to the 1990s [7], however, the advent of LLM-driven multi-agent ecosystems [8] since 2023 has made “ Agentic AI ” a distinctly new and rapidly expanding research focus [5]. Traditional agents primarily automated narrow tasks, such as information retrieval [9, 10], data summarization [11], and dialogue response [12]. Typically driven by single step logic [13] or scripted rules, these agents lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem solving [14, 15].
In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), enabled by LLMs and tool use [16]. These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines [17, 18, 19]. This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent and decentralized behavior.
@@ 1.2 Trust, Risk, and Security Challenges in Agentic AI
The rise of Agentic AI introduces complex challenges that extend beyond those of traditional ML systems. Unlike deterministic agents [20], LLM-based multi-agent systems exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors. These characteristics necessitate a robust approach to have trust and ensuring trustworthiness, in high-stakes domains such as healthcare, finance, and law.
Trust and Trustworthiness
@@
In the context of Agentic AI, trust refers to a user willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner, thereby deserving that trust. This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications. The stochastic nature of LLM reasoning [21] further complicates trustworthiness, as randomness in output generation, while promoting creativity and ensemble diversity, introduces inconsistency and non-repeatability. These factors hinder traceability, verification, and compliance auditing, all of which are essential for trustworthy systems.
Security Risks in Multi-Agent Systems
@@
Security risks escalate as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [22]. For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents but inadvertently leak sensitive data or violate compliance protocols (e.g., GDPR or HIPAA) if safeguards are inadequate. The distributed and cooperative nature of multi-agent systems introduces unique vulnerabilities that traditional security measures often fail to address.
Specific Threats: Prompt Injection, Spoofing, and Impersonation
@@
LLM-based agents are susceptible to novel security threats exploiting their language-based interfaces and cooperative behaviors. One prominent threat is prompt injection [23], where attackers craft inputs with hidden or malicious instructions to manipulate agent behavior. Recent studies [24] describe “prompt infection,” a phenomenon where malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [25]. Such corrupted outputs can lead to cascading effects, including data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
@@
Another critical vulnerability involves spoofing and impersonation [26]. Spoofing entails faking an entity’s identity or credentials (e.g., using a false name or address) to deceive a system, while impersonation involves actively pretending to be a specific agent or user to gain trust, access, or privileges. In multi-agent systems, where agents rely on credentials or tokens for authentication during coordination, an adversary could steal an API key or mimic a trusted peer to issue unauthorized commands or access sensitive information.
Privacy Risks in Multi-Agent LLM Systems
@@
LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks [27]. In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate.
Limitations of Traditional Frameworks
Traditional evaluation and safety frameworks, designed for deterministic agents or conventional AI systems, are increasingly inadequate for addressing the autonomy and complexity of Agentic AI. These frameworks often fail to account for emergent behaviors, inter-agent dependencies, or the dynamic nature of LLM-driven systems, leaving gaps in security, trust, and risk management.
Towards a TRiSM Framework for Agentic AI
To address these gaps, recent frameworks like AI TRiSM (Trust, Risk, and Security Management) [28, 29] propose lifecycle-level controls, including explainability, secure model orchestration, and privacy management. These components are vital for deploying Agentic AI in high-stakes domains such as finance, healthcare, and defense, where trust, security, and accountability are non-negotiable.
@@ 1.3 Problem Statement
Agentic AI systems increasingly rely on the coordination of multiple LLM-powered agents to perform complex, long-horizon tasks. However, no unified framework exists for managing their Trust, Risk, and Security Management (AI TRiSM) across dynamic, multi-agent workflows. Recent failures, such as, an autonomous research agent citing unverifiable medical sources or collaborative swarms generating inconsistent forecasts due to memory drift, highlight that traditional safeguards are inadequate [30]. These traditional systems are vulnerable to orchestration breakdowns, agent collusion, data leakage, and opaque decision-making pathways. To ensure safety and reliability in Agentic AI systems, a structured TRiSM perspective is urgently needed. This study summarizes emerging findings and offers actionable guidelines to build secure, transparent, and trustworthy Agentic AI ecosystems.
@@ 1.4 Scope and Objectives
Objectives This study is a review article that maps the TRiSM (Trust, Risk, and Security Management) [28] landscape as it applies to LLM-powered, multi-agent “Agentic AI” systems. We structure the discussion around four foundational pillars of TRiSM [31]: explainability, security, lifecycle governance, and privacy. We argue that conventional safeguards developed for bounded, rule-based agents are inadequate for managing the complexities of open-ended, self-directed agent collectives. The objective of this review is to offer researchers and practitioners a coherent and actionable reference for the design, deployment, and governance of trustworthy Agentic AI ecosystems.
For the sake of this study, we define Agentic AI as Multi-Agent Systems (MAS) [32] powered by LLMs that exhibit autonomous planning, tool use, and memory retention, with or without human oversight. In contrast to classical MAS, these systems exhibit emergent reasoning capabilities and dynamically adapt their workflows based on interactions with tools, users, and other agents.
Scope This paper is intended as a conceptual survey and synthesis. It does not propose a new Agentic AI architecture or introduce synthetic benchmark evaluations. Rather, it integrates emerging perspectives on trust, risk, and security to support future research and responsible development in this evolving area.
Necessity of this Survey Agentic AI research has largely centered on agent modeling, planning, and task coordination, while comparatively overlooking critical dimensions such as adversarial robustness, lifecycle governance, and system-level explainability. As these systems are increasingly deployed in high-stakes domains - such as healthcare, scientific research, and finance—the absence of an integrated TRiSM framework exposes stakeholders to opaque decision-making and unmanaged risks. This survey addresses that gap by synthesizing TRiSM principles in the context of LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policymakers.
@@ 1.5 Contributions
  1. 1.
    We propose the first TRiSM framework that is explicitly tailored to systems in which several LLM-powered agents collaborate and cooperate. The framework is organized around four pillars: Explainability, Lifecycle Governance, Application Security, and Model Privacy. It targets the specific governance challenges created by autonomous decision making, inter-agent coordination, and persistent memory (see Fig. 1 and Sec. 5).
  2. 2.
    We offer a systematic taxonomy of risks that emerge when multiple agents interact (Sec. 6). This taxonomy discusses prompt injection, memory poisoning, collusive failure, emergent misbehavior, and related threats, and it shows why defenses designed for single-model LLM applications are not sufficient.
  3. 3.
    We introduce a five-dimension evaluation framework that covers Trustworthiness, Explainability, User-centered Performance, Coordination, and Composite Assessment (Sec. 7.3). To support this template, we define two new metrics: the Component Synergy Score (CSS), which measures the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which measures how correctly and efficiently agents invoke external tools.
  4. 4.
    We review state-of-the-art techniques for each TRiSM pillar (Secs. 78) and examine methods such as Local Interpretable Model-Agnostic Explanations (LIME), SHapley Additive exPlanations (SHAP), decision provenance, prompt hygiene, sandboxing, and ModelOps (model operations) pipelines. These techniques must be critically re-evaluated and adapted to suit environments involving multiple interacting LLM agents.
  5. 5.
@@
1.6 Paper Organization
The remainder of this paper is structured as follows. Section 2 discuss the related work. Section 3 presents the literature search methodology. Section 4 reviews the core characteristics and architectures of Agentic AI systems, providing a foundation for the TRiSM analysis. Section 5 introduces the TRiSM framework and its four pillars: Explainability, ModelOps, Application Security, and Model Privacy. Section 6 outlines key risks, vulnerabilities, and attack surfaces in multi-agent AI ecosystems. Section 7 discusses trust-building and explainability techniques, while Section 8 covers security and privacy-enhancing methods. Section 9 synthesizes insights and proposes directions for research and policy. Section 10 concludes the paper with final reflections on the implications of TRiSM for the responsible development of Agentic AI. A list of key terms used throughout this paper is provided in Table A.1.
@@ 2 Related Work
2.1 LLM-Based Multi-Agent Surveys (Technical Focus)
Prior surveys on LLM-driven multi-agent systems predominantly focus on technical dimensions such as system architectures, agent capabilities, and domain-specific applications of Agentic AI, yet they rarely address critical TRiSM (Trust, Risk, and Security Management) considerations. For example, reviews like [33] and [34] delve into simulated environments, inter-agent communication protocols, and performance benchmarks. These works excel at detailing how agents coordinate or scale but fall short in examining adversarial threats, such as prompt injection attacks or agent collusion, lifecycle governance (e.g., continuous monitoring or regulatory compliance), or system-wide explainability, especially for emergent behaviors in complex systems.
More specialized surveys, such as [35] on natural language communication and [36] on collaboration mechanisms, explore interaction protocols and practical use cases in domains like question-answering systems, 5G networks, and Industry 5.0. While [35] briefly notes security risks tied to communication channels, and [36] mentions deployment contexts, both remain narrowly focused on coordination strategies. They lack deeper analysis of robustness against adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, or explainability challenges at scale.
2.2 Trustworthy/Responsible AI Surveys (Broad Trust Focus)
Surveys addressing TRiSM in AI, such as [38], tackle ethical and risk-related issues including alignment with human values, fairness in decision-making, and defenses against privacy attacks (e.g., data leakage). However, these works typically target general ML systems [39] and overlook the unique dynamics of LLM-based multi-agent systems. For instance, while they may advocate for privacy-preserving techniques like differential privacy, they do not address privacy risks in inter-agent data sharing or the distributed nature of LLM-driven systems. Moreover, critical TRiSM aspects [40] like governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents) are often absent. As a result, their recommendations, while valuable for single-agent or centralized AI, do not fully translate to the complexities of multi-agent coordination in high-stakes domains.
2.3 Current Study (LLM Multi-Agent)
This review distinguishes itself by weaving TRiSM principles: adversarial robustness, end-to-end lifecycle governance, and system-wide explainability, into the study of LLM-based multi-agent systems, with a specific lens on high-stakes domains like healthcare, science, and finance. Unlike prior studies (see Table 1), which offer limited insight into governance and security, this work positions these elements as foundational design priorities. For example, it addresses targeted threats like backdoor attacks in distributed LLMs and governance challenges such as ensuring compliance with regulations (e.g., HIPAA in healthcare or GDPR in finance) across autonomous agents - areas largely ignored by technical surveys. This work provides actionable guidance for researchers, engineers, and policymakers, and fills critical gaps in both the technical multi-agent literature and broader responsible AI landscape.

### v3

@@ Abstract
Agentic AI systems, built upon large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligence, autonomy, collaboration, and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based Agentic Multi-Agent Systems (AMAS). We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents. We then adapt and extend the AI TRiSM framework for Agentic AI, structured around four key pillars: Explainability, ModelOps, Security, Privacy and Governance, each contextualized to the challenges of multi-agent LLM systems. A novel risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation. To support practical assessment in Agentic AI works, we introduce two novel metrics: the Component Synergy Score (CSS), which quantifies the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which evaluates the efficiency of tool use within agent workflows. We further discuss strategies for improving explainability in Agentic AI , as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance. The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, outlining critical directions to align emerging systems with TRiSM principles for safe, transparent, and accountable operation.
@@ 1.1 Background and Motivation
The global market for Artificial Intelligence (AI) agents is projected to grow from \$5.4 billion in 2024 to \$7.6 billion in 2025  [1]. By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [2]. An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals  [3], has undergone a fundamental transformation. While early agents were task-specific and deterministic, modern systems have evolved significantly ,increasing operational autonomy by over 70% on benchmark tasks [4]. These Agentic AI systems now integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [5].
This rapid evolution presents new opportunities but also brings challenges in coordination, control, transparency, and risk assessment across heterogeneous agents. In particular, as the shift toward agentic AI architectures amplifies and evolves, longstanding concerns around safety, fairness, interpretability, and systemic harms resurface. While such issues have been central to machine learning (ML) systems broadly, they become more promiment in Agentic AI, where autonomous agents operate in high-stakes domains like healthcare, law, and public services, areas where their decisions can carry profound and lasting societal consequences.
From Isolated Agents to Agentic AI The autonomous software agents date back to the 1990s [6], however, the advent of LLM-driven multi-agent ecosystems [7] since 2023 has made “Agentic AI”, a distinctly new and rapidly expanding research focus [5]. Traditional agents primarily focussed on automating limited tasks, such as information retrieval [8], data summarization [9], and dialogue response [10]. Typically driven by single step logic [11] or scripted rules, these traditional agents lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem solving .
In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), and enabled by LLMs and tool use [12]. These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines [13]. This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent and decentralized behavior.
@@ 1.2 Trust, Risk, and Security Challenges in Agentic AI
The rise of Agentic AI introduces new challenges that extend beyond those of traditional ML systems. For example, unlike deterministic agents [14], LLM-based multi-agent systems exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors. These characteristics necessitate a robust approach to have trust and ensuring trustworthiness, in high-stakes domains such as healthcare, finance, and law. Below, we discuss some of these challanges:
  1. 1.
@@
    Trust and Trustworthiness In the context of Agentic AI, trust refers to a user willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner, thereby deserving that trust. This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications. The stochastic nature of LLM reasoning [15] further complicates trustworthiness, as randomness in output generation, while promoting creativity and ensemble diversity, introduces inconsistency and non-repeatability. These factors hinder traceability, verification, and compliance auditing, all of which are essential for trustworthy systems.
  2. 2.
@@
    Security Risks in Multi-Agent Systems Security risks escalate as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [16]. For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents but inadvertently leak sensitive data or violate compliance protocols (e.g., GDPR or HIPAA) if safeguards are inadequate. The distributed and cooperative nature of multi-agent systems introduces unique vulnerabilities that traditional security measures often fail to address.
  3. 3.
@@
    Specific Threats: Prompt Injection, Spoofing, and Impersonation LLM-based agents are susceptible to novel security threats exploiting their language-based interfaces and cooperative behaviors. One prominent threat is prompt injection [17], where attackers design inputs with hidden or malicious instructions to manipulate agent behavior. Recent studies [18] describe “prompt infection” as a phenomenon where malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [19]. Such corrupted outputs can lead to cascading effects, including data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
@@
    Another critical vulnerability involves spoofing and impersonation [20]. Spoofing means faking an entity’s identity or credentials (e.g., using a false name or address) to deceive a system, while impersonation involves actively pretending to be a specific agent or user to gain trust, access, or privileges. In multi-agent systems, where agents rely on credentials or tokens for authentication during coordination, an adversary could steal an API key or mimic a trusted peer to issue unauthorized commands or access sensitive information.
  4. 4.
@@
    Privacy Risks in Multi-Agent LLM Systems LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks [21]. In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate.
Traditional evaluation and safety frameworks, designed for deterministic agents AI systems, are inadequate for addressing the autonomy and control of Agentic AI. To address these gaps, frameworks like AI TRiSM (Artificial Intelligence Trust, Risk, and Security Management) [22, 23] proposes lifecycle-level controls, including explainability, secure model orchestration, and privacy management. AI TRiSM is a framework to help organizations build AI systems that people can trust. The goal is to focus not just on technical performance, but also on ethical values like fairness, transparency, and security. The framework brings together practices that support governance, strengthen reliability and robustness, and protect data and make sure that AI systems are not only effective, but also aligned with human values.
@@ 1.3 Problem Statement
Agentic AI systems increasingly rely on the coordination of multiple LLM-powered agents to perform complex, long-horizon tasks. However, no unified framework exists for managing the trust, risk, and security across dynamic as well as multi-agent workflows. Recent accidents, such as, an autonomous research agent citing unverifiable medical sources or collaborative swarms generating inconsistent forecasts due to memory drift, highlight that traditional safeguards are inadequate [24]. These traditional systems are vulnerable to orchestration breakdowns, agent collusion, data leakage, and opaque decision-making pathways. To ensure safety and reliability in Agentic AI systems, a structured TRiSM perspective is therefore urgently needed.
@@ 1.4 Scope and Objectives
Objectives This study is a review article that maps the TRiSM (Trust, Risk, and Security Management) [22] framework to LLM-powered, multi-agent “Agentic AI” systems. We structure the discussion around four foundational pillars of TRiSM [25]: explainability, security, lifecycle governance, and privacy. The objective of this review is to offer researchers and practitioners a coherent and actionable reference for the design, deployment, and governance of trustworthy Agentic AI ecosystems.
For the purposes of this study, we define Agentic AI as Multi-Agent Systems (MAS) [26] powered by LLMs that exhibit autonomous planning, tool use, memory retention, and emergent reasoning capabilities, with or without human supervision. These LLM-based agents can dynamically adapt workflows based on interactions with tools, users, and other agents, distinguishing them from traditional symbolic or rule-based agents.
Scope This paper is a conceptual survey and synthesis. It does not propose a new architecture for Agentic AI, nor does it include synthetic benchmark evaluations. Instead, it integrates current perspectives on AI TRiSM to inform future research and the responsible development of Agentic AI systems. The scope of this paper is confined to LLM based Agentic AI systems and AMAS  [^3] Throughout this paper, AMAS designates multi-agent systems built on LLMs.
Necessity of this Survey Agentic AI research has largely centered on agent modeling, planning, and task coordination, while comparatively overlooking critical dimensions such as adversarial robustness, lifecycle governance, and system-level explainability. As these systems are increasingly deployed in high-stakes domains - such as healthcare, scientific research, and finance the absence of an integrated TRiSM framework exposes stakeholders to opaque decision-making and unmanaged risks. This survey addresses that gap by synthesizing TRiSM principles in the context of LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policymakers.
@@ 1.5 Contributions
  • We propose the first TRiSM framework for Agentic AI systems. This framework is explicitly tailored to systems in which several LLM-powered agents collaborate and cooperate. The framework is organized around pillars: Explainability, ModelOps, Security, Privacy and Governance around autonomous decision to guide the development and use of safe and ethical Agentic AI systems. (see Fig. 1 and Sec. 6).
  • We offer a systematic taxonomy of risks that emerge when multiple agents interact (Sec. 5). This taxonomy discusses prompt injection, memory poisoning, collusive failure, emergent misbehavior, and related threats commonly seen in AMAS.
  • We introduce an evaluation dimensions for Agentic AI systems that covers Trustworthiness, Explainability, User-centered Performance, Coordination, and Composite Assessment (Sec. 8). To support this template, we define two new metrics: the Component Synergy Score (CSS), which measures the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which measures how correctly and efficiently agents invoke external tools. Our aim is to achieve high performance in both dimensions, accuracy and usability, without sacrificing one.
  • We contribute to the management and control of Agentic AI by mapping existing TRiSM techniques such as explainability methods, prompt hygiene, decision provenance, sandboxing, and ModelOps pipelines, to AMAS environments. We critically analyze their limitations and adaptation needs, providing a foundation for re-engineering these techniques to fit the demands of interacting LLM agents (Secs. 79).
@@ 2 Related Work
Xi et al. (2025) [33] ✓ (single-/multi-agent,
human-AI coop)
~ (open problems)
Luo et al. (2025) [34] ~ ~ ✓ (science, games, etc.) ~ (open challenges)
Zou et al. (2025) [35] ~ ✓ (e.g., finance,
healthcare)
~ (challenges & opp.)
Wang et al. (2025) [36] ~ ~ ~ (healthcare domain) ~ (future opportunities)
@@
LLM-Based Multi-Agent Surveys (Technical Focus)
Prior surveys on LLM-driven AMAS predominantly focus on technical dimensions such as system architectures, agent capabilities, and domain-specific applications of Agentic AI. However, there remains a gap in addressing governance and policy aspects. For example, reviews like [27] and [28] delve into simulated environments, inter-agent communication protocols, and performance benchmarks. These works mainly detail how agents coordinate or scale but fall short in examining adversarial threats, such as prompt injection attacks or agent collusion, lifecycle governance (e.g., continuous monitoring or regulatory compliance), or system-wide explainability, especially for emergent behaviors in complex systems.
More specialized surveys, such as [29], focus on natural language communication, while [30] details collaboration mechanisms, exploring interaction protocols and practical use cases in domains like question-answering systems, 5G networks, and Industry 5.0. Although [29] highlights security risks associated with communication channels, and [30] discusses deployment contexts, both remain narrowly centered on coordination strategies. There is a need for deeper analysis of robustness against adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, and explainability challenges at scale.
Trustworthy/Responsible AI Surveys (Broad Trust Focus)
There are some review articles that address TRiSM in AI, such as [32], tackle ethical and risk-related issues including alignment with human values, discuss fairness in decision-making [37], and defenses against privacy attacks (e.g., data leakage). However, these works typically target general ML systems and overlook the unique dynamics of LLM-based multi-agent systems. For instance, while they may advocate for privacy-preserving techniques like differential privacy, they do not address privacy risks in inter-agent data sharing or the distributed nature of LLM-driven systems. Moreover, critical TRiSM aspects [38] like governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents) are often absent in recent discussions.
Current Study (Multi-Agent Systems and LLMs)
This review distinguishes itself by mapping TRiSM principles into the study of AMAS. Unlike prior studies (see Table 1), which offer limited insight into governance and policy, this work operationalizes TRiSM principles in practice. It provides actionable guidance for researchers, engineers, and policymakers, addressing critical gaps in both the technical AMAS literature and the broader responsible AI landscape.